# Bridging Entanglement and Magic Resources through Operator Space

### v1

@@ Abstract
Introduction.— The growth of non-classical resources in quantum dynamics indicates a necessary piece of the puzzle separating classical and quantum simulability. Understanding this separation is essential to a complete characterization of such systems, both from an algorithmic and from a physical perspective. Arguably the most famous of these resources is entanglement: it both governs the efficiency of tensor network methods Verstraete and Cirac (2006); Schuch et al. (2008), and witnessing quantum critical phase transitions Vidal et al. (2003); Calabrese and Cardy (2004); Hastings (2007); Eisert et al. (2010); Silvi et al. (2010). Similarly, the concept of non-stabilizerness—or magic—has emerged as another key resource of non-classicality. It is well-known that Clifford dynamics can be simulated efficiently on a classical computer by tracking the $N$ stabilizer generators of an $N$-qubit state Gottesman (1998); Aaronson and Gottesman (2004). Including a $T$-gate in the gateset is all that is required to remove these generators and achieve universality—and indeed the best known Clifford+$T$ simulators scale exponentially in the number of $T$-gates Pashayan et al. (2022); Bravyi et al. (2019). This departure from stabilizerness can be formalized rigorously in the theory of magic, where $T$-count forms just one example aspect Howard and Campbell (2017); Beverland et al. (2020); Leone et al. (2022); Liu and Winter (2022). Recently, magic resources have also been found to play a key role in phase transitions White et al. (2021); Niroula et al. (2024); Catalano et al. (2024); Fux et al. (2024a), dynamical complexity Leone et al. (2022); Garcia et al. (2023); Oliviero et al. (2024); Gu et al. (2024); Ahmadi and Greplova (2024); Odavic et al. (2024a), and (pseudo-)randomness Haferkamp et al. (2022a); Leone et al. (2021).
@@
Given the clear role each of these concepts individually play in both quantum information theory and many-body physics, one would hope to transitively understand the direct connection between the two. A question of foundational importance is therefore: How are the two disparate resources of magic and entanglement related to one other? The solution is not a priori obvious, at least in the usual quantum state setting. For instance, entanglement tends to grow maximally in Clifford circuits. On the other hand, product states can have high magic resource. And although details of the entanglement spectrum Chamon et al. (2014) can serve as witness to magic in a state Tirrito et al. (2024a); Turkeshi et al. (2023), even a single $T$-gate can make this spectrum consistent with random matrix theory predictions Zhou et al. (2020). Even more puzzlingly, there appears no clear connection to the integrability of the underlying dynamics in a many-body setting. Locally-interacting integrable and chaotic models alike produce linearly growing entanglement after a quench Calabrese and Cardy (2005); Chiara et al. (2006); Prosen and Znidarič (2007); Mezei and Stanford (2017); Nahum et al. (2017), while recent evidence suggests that state magic resources saturate in logarithmic time Turkeshi et al. (2024); Tirrito et al. (2024b); Odavic et al. (2024b). Despite this, there has been recent interest in studying the interplay of, and boundary between, these fundamental resources Masot-Llima and Garcia-Saez (2024); Fux et al. (2024b); Gu et al. (2024); Frau et al. (2024).
@@
Figure 1: Depiction of of key quantities investigated in this work. (a) The local-operator entanglement (LOE) is the entanglement $E_{A}^{(\alpha)}$ of the Choi state of an operator, $|O_{U}\rangle\!\rangle$, across some (doubled-space) spatial bipartition $A:\bar{A}$ for quantum $\alpha$-Rényi entanglement entropies. (b) The extensive scaling of this quantity witnesses chaos, both in the sense of information scrambling and non-integrability. (c) In addition to being sensitive to entanglement, we show that the LOE equally probes magic resources and operator stabilizer entropies, unifying the three corresponding simulation techniques under the one umbrella.
@@
Here, we pose a new understanding of this relationship. We show that entanglement and magic are conceptually and quantitatively connected. We are able to do this by a shift in perspective from states to operators, noticing that free stabilizer resources are product in operator space. More specifically, consider an operator $O$ in the Heisenberg picture with respect to some evolution $U$, i.e., $O_{U}=U^{\dagger}OU$. By the Choi–Jamiołkowski isomorphism (CJI) $O_{U}$ has a dual pure state, $|O_{U}\rangle\!\rangle:=(O_{U}\otimes\mathds{1})\ket{\phi^{+}}$, where |$\ket{\phi^{+}}$ is the (normalized) maximally entangled state on a doubled Hilbert space. Entanglement and magic then have ready generalizations to operators through $|O_{U}\rangle\!\rangle$. The entanglement $E_{A}^{(\alpha)}(O_{U})$ of this state—once appropriately arranged into spatial partitions—is termed the local-operator entanglement (LOE) Prosen and Pižorn (2007), and has been studied in a variety of many-body systems Prosen and Znidarič (2007); Prosen and Pižorn (2007); Pižorn and Prosen (2009); Dubail (2017); Jonay et al. (2018); Alba et al. (2019); Bertini et al. (2020a, b); Alba (2021); see Fig. 1.
@@
The above relation Eq. (1) is the subject of the remainder of this work, with the full technical version to be found towards the end of the manuscript (Thm. 2). Our result sits in stark contrast to the corresponding state resource theories, where no such relation exists: states with high magic can be unentangled, while zero-magic (stabilizer) states can have maximal entanglement. Beyond the foundational significance of Eq. (1), there are operational implications. Chaotic dynamical systems have linearly scaling LOE with time, whereas it grows at-fastest logarithmically for integrable dynamics Prosen and Znidarič (2007); Prosen and Pižorn (2007); Pižorn and Prosen (2009); Dubail (2017); Jonay et al. (2018); Alba et al. (2019); Bertini et al. (2020a, b); Alba (2021). Our main result Eq. (1) additionally means that an extensive growth of magic resources is necessary ingredient for quantum chaos (see Cor. 3). Interpreting this result, chaotic growth of LOE implies the impossibility of efficient simulation of those dynamics using standard: tensor network, stabilizer, and Pauli truncation methods (see Cor. 4). LOE therefore serves as a unifying signature of non-classicality.
@@
Average Operator Entanglement from Magic.— Before detailing the full version of Eq. (1) (Thm. 2), we first review the pertinent monotones (summarized in Table 1) and introduce the relevant unitary ensembles (depicted in Fig. 2). Our many-body system of interest $\mathcal{H}$ is an $N$-qubit Hilbert space with total dimension $D=2^{N}$, and denote by $\mathcal{U}_{N}$, $\mathcal{C}_{N}$, and $\mathcal{P}_{N}$ the $N$-qubit unitary, Clifford and Pauli groups respectively. By $U\sim\mathcal{U}_{N}$, we mean that $U$ is sampled uniformly according to the unitarily invariant (Haar) measure on the $N$-qubit unitary group, with an equivalent expression for the Cliffords.
Consider a Heisenberg operator $O_{U}=U^{\dagger}OU$ for some non-trivial initial Pauli operator $O\in\mathcal{P}_{N}\backslash\{\mathds{1}\}$ and a unitary propagator $U$. Through the CJI, the LOE is defined as the bipartite entanglement of the Choi state $|O_{U}\rangle\!\rangle$ across a chosen bipartition $\mathcal{H}=\mathcal{H}_{A}\otimes\mathcal{H}_{\bar{A}}$, $E^{(\alpha)}_{A}(O_{U}):=S^{(\alpha)}(\operatorname{tr}_{\bar{A}}[|O_{U}\rangle\!\rangle\!\langle\!\langle O_{U}|])$, where $S^{(\alpha)}$ represents the (quantum) $\alpha-$Rényi entropy, $S^{(\alpha)}(\rho):=({1-\alpha})^{-1}\log(\operatorname{tr}[\rho^{\alpha}])$. Through abuse of notation, $\operatorname{tr}_{\bar{A}}$ refers to a partial trace over subsystem ${\bar{A}}$ in the doubled space $\mathcal{H}_{\bar{A}}\otimes\mathcal{H}_{\bar{A}}$ with dimension $D_{\bar{A}}^{2}$. We will leave unspecified the exact choice of (spatial) bipartition and Rényi index $\alpha$, unless otherwise stated. When determining the average value of the LOE for classes of circuits, it will also be convenient to study the operator purity of $O_{U}$, defined in the usual way through
$E^{(\text{pur})}_{A}(O_{U}):=\exp(-E^{(2)}_{A}(O_{U})).$ (2)
@@
To first provide intuition on the interplay between LOE and magic resources, consider Clifford evolution, $C\in\mathcal{C}_{N}$. It is immediate to see that for any initial Pauli operator $O$ will evolve to some other Pauli operator from the definition of the Clifford group and so the LOE is preserved to be zero. We again stress that no such relation between entanglement and magic holds for states. This invites the question of whether a more quantitative relation can be derived: if a unitary has only a few non-Clifford gates, how does the LOE grow? We will first answer this question by computing the average LOE for ensembles with a tunable magic monotone and show that there is a precise dependence.
@@
Figure 2: Any unitary can be decomposed as a global Clifford unitaries, $C_{0},C_{1}\in\mathcal{C}_{N}$, sandwiching a non-Clifford unitary on only $\nu(U)/2$ qubits, $V\in\mathcal{U}_{\nu}$ (top), where $\nu(U)$ is the unitary nullity Jiang and Wang (2023). A unitary with $T$-count $\tau(U)$ can be decomposed as layers of $N$-qubit Clifford unitaries $C_{i}\in\mathcal{C}_{N}$ interspersed with $\tau(U)$ single-site $T$-gates (bottom). The $\nu$-compressible and $T$-doped ensembles are constructed from uniformly sampling $\{C_{0},C_{1},V\}$ and $\{C_{i}\}_{0\leq i\leq\tau}$ over the Clifford and unitary groups as appropriate, and are denoted $\mu_{\nu}$ and $\mu_{\tau}$ respectively.
@@
Consider deep Clifford circuits doped with some non-Clifford gates. That is, random Clifford circuits interspersed by $\tau(U)$ single-site $T$-gates, where $\tau(U)$ is the familiar $T$-count magic monotone (see Table 1). If we randomize the Clifford components, such an ensemble of unitaries $\mu_{\tau}$ is called the $T$-doped Clifford ensemble Haferkamp et al. (2022a); Leone et al. (2021) (see Table 1 and Fig. 2). Harnessing Weingarten techniques for averaging over the Clifford group Zhu et al. (2016); Roth et al. (2018); Leone et al. (2021), we can readily compute the operator purity for this example,
$\int_{U\sim\mu_{\tau}}\!\!\!\!\!\!{E^{(\text{pur})}_{A}(O_{U})}=\exp\left(-\log(\tfrac{4}{3})\tau+\mathcal{O}\left(\frac{({4}/{3})^{\tau}}{D}\right)\right),$ (3)
where we choose a bipartition of size $N_{A}=N/2$ for the computation of the LOE. The full details of the proof of Eq. (3) (and the below Eq. (6)) can be found in Appendix B, together with exact expressions for arbitrary $\tau(U)$, $N$, and $N_{A}$. Interestingly, the dependence on $(3/4)^{\tau}$ in Eq. (3) is reminiscent of that observed for the linear stabilizer Rényi entropy of the $T-$doped ensemble for large $N$ Leone et al. (2022); Haug et al. (2024). Key for our goals, Eq. (3) shows a linear relationship between $\tau(U)$ and entanglement (recalling Eq. (2)). This result indicates already a proportionality of operator entanglement with magic resources, at least in the typical (average) case. However, while useful practically, it is not always possible to determine the exact $T$-count of a unitary.
@@
$T$-count Minimum number $\tau(U)$ such that $U=\prod_{i}^{\tau}C_{i}T$, for $C_{i}\in\mathcal{C}_{N}$.
@@
A more concrete magic resource stemming from the algebraic structure of the Clifford group, is the stabilizer nullity Beverland et al. (2020). For quantum states |$\ket{\psi}$, this is defined in terms of the cardinality of its stabilizer group: the number of Pauli operators $P$ which satisfy $|\langle\psi|P|\psi\rangle|=1$. The nullity of a unitary operator can be defined analogously Jiang and Wang (2023),
$\nu(U):=2N-\log_{2}(|\mathrm{Stab}(\ket{U})|).$ (4)
Here, $\mathrm{Stab}(\ket{U})$ is the stabilizer group of the Choi state |$\ket{U}$, equal to $\mathrm{Stab}(\ket{U})=\{P_{1},P_{2}\in\mathcal{P}_{N}:\operatorname{tr}[P_{1}U^{\dagger}P_{2}U]/D=\pm 1\}$. The unitary nullity always takes a non-negative integer value, upper-bounds the largest state nullity of $U$ acting on a stabilizer state, and can be shown to lower-bound the $T$-count Jiang and Wang (2023). Moreover, it gives rise to a powerful representation theorem for any unitary which we present below, building upon previous compressibility results Leone et al. (2024a, b); Gu et al. (2024).
Proposition 1.
Any unitary $U$ with nullity $\nu$ can be decomposed as
$U=C_{0}(V_{\nu/2}\otimes\mathds{1}_{N-\nu/2})C_{1},$ (5)
where $C_{0},C_{1}\in\mathcal{C}_{N}$ can act globally, while $V_{\nu/2}\in\mathcal{U}_{\lceil\nu/2\rceil}$ acts on exactly $\lceil\nu(U)/2\rceil$ qubits.
@@
The proof for this follows from the fact that the space spanned by the Paulis in the complement of $\mathrm{Stab}(U)$ (of size $2^{\nu}$), together with their image, can both be mapped to a local space of only $\nu/2$ qubits through pre- and post-processing using Cliffords. Moreover, $P\in\mathrm{Stab}(U)$ can be mapped through the action of another Clifford, which commutes with the non-Clifford part $V$—see App. A. ∎
In contrast to classifying magic in terms of $T$-count, any unitary can be decomposed according to Eq. (5). Of course, for the majority of (high-complexity) unitaries in $\mathcal{U}_{N}$, $\nu(U)=2N$ and so this representation is trivial Poulin et al. (2011). If $\nu(U)<2N$, we call such a unitary $\nu$-compressible. In comparison to similar results presented in Refs. Leone et al. (2024a, b); Gu et al. (2024), Prop. 1 is an improvement through halving the number of qubits in the compression. In particular, Prop. 1 together with a result from Ref. Jiang and Wang (2023) implies that a Clifford circuit with $\tau(U)$ $T$-gates can be compressed to a non-Clifford $V$ acting on at-most $\lceil\tau(U)/2\rceil$ qubits; c.f. Thm. 2 from Ref. Leone et al. (2024b). It is an interesting question to further investigate the implications of this improvement.
Returning to the question of the LOE’s sensitivity to magic resources, we next compute the average-case operator purity for a unitary with a given unitary nullity $\nu(U)$. We define the $\nu$-compressible ensemble of unitaries $\mu_{\nu}$ as that generated from independently randomly sampling $C_{0},C_{1}\in\mathcal{C}_{N}$ and $V\in\mathcal{U}_{\nu/2}$ in Eq. (5) uniformly over the Clifford and unitary groups respectively. Such a unitary ensemble is of indpendent interest, and should find application in other areas of many-body physics. The average operator purity of an initial Pauli $O\in\mathcal{P}_{N}\backslash\{\mathds{1}\}$ for $\nu$-compressible unitaries is
$\int_{U\sim\mu_{\nu}}\!\!\!{E^{(\text{pur})}_{A}(O_{U})}=\exp\left(\!-\nu+2+\mathcal{O}\left(\frac{1}{4^{\nu}}+\frac{2^{\nu}}{D}\right)\right),$ (6)
where we use Weingarten techniques for taking Haar averages Collins and Sniady (2006); Zhu et al. (2016); Roth et al. (2018); Mele (2024). Again, via Eq. (2) we observe a linear dependence of operator entanglement with magic monotone $\nu(U)$; c.f. Eq. (1).
Before moving on to an exact bound between LOE and magic resources (without averaging), we would like to compare Eqs. (3) and (6). The ensembles $\mu_{\tau}$ and $\mu_{\nu}$ coincide only for two cases: $\tau=\nu=0$, in which case they correspond uniform measure over Clifford group, and $\tau\to\infty$, $\nu\to 2N$, in which case they correspond to the unitary Haar ensemble. In the former case, as previously discussed, the LOE is trivially equal to zero. In the latter case, we can use our previous results to arrive at a Page-curve-like expression for the average LOE from Haar random dynamics,
@@
Note that this result generalizes one given in Ref. Kudler-Flam et al. (2021) to beyond only leading order in $1/D$.
Unifying Entanglement and Magic.— We will now show that exact bounds can be derived between these two quantities for arbitrary Rényi indices and any given evolution, resulting in Eq. (1). In order to prove our main result, we first review one further magic monotone. The OSE is defined as the entropy of the distribution of square amplitudes of the Heisenberg operator $O_{U}$ written in the Pauli basis Dowling et al. (2024),
$M^{(\alpha)}(O_{U}):=(1-\alpha)^{-1}\log\sum_{P\in\mathcal{P}}\left(D^{-1}\operatorname{tr}[O_{U}P]\right)^{2\alpha}.$ (8)
This quantity generalizes a popular state measure of magic Leone et al. (2022); Haug and Piroli (2023); Leone and Bittel (2024), lower-bounding other magic monotones while satisfying a light-cone bound for local dynamics, and (most pertinent to our results) its scaling dictates the efficiency of Pauli-truncation methods. Using the OSE, together with Prop. 1 and the average-case computations Eqs. (3) and (6), we are now in a position to state our main result (Eq. (1)) in full technical detail.

### v2

@@ Abstract
Introduction.— The growth of non-classical resources in quantum dynamics indicates a necessary piece of the puzzle separating classical and quantum simulability. Understanding this separation is essential to a complete characterization of such systems, both from an algorithmic and from a physical perspective. Arguably the most famous of these resources is entanglement: it both governs the efficiency of tensor network methods Verstraete and Cirac (2006); Schuch et al. (2008), and witnesses quantum critical phase transitions Vidal et al. (2003); Calabrese and Cardy (2004); Hastings (2007); Eisert et al. (2010); Silvi et al. (2010). Similarly, the concept of non-stabilizerness—or magic—has emerged as another key resource of non-classicality. It is well-known that Clifford dynamics can be simulated efficiently on a classical computer by tracking the $N$ stabilizer generators of an $N$-qubit state Gottesman (1998); Aaronson and Gottesman (2004). Including a $T$-gate in the gateset is all that is required to remove these generators and achieve universality—and indeed the best known Clifford+$T$ simulators scale exponentially in the number of $T$-gates Pashayan et al. (2022); Bravyi et al. (2019). This departure from stabilizerness can be formalized rigorously in the theory of magic, where $T$-count forms just one example aspect Howard and Campbell (2017); Beverland et al. (2020); Leone et al. (2022); Liu and Winter (2022). Recently, magic resources have also been found to play a key role in phase transitions White et al. (2021); Niroula et al. (2024); Catalano et al. (2024); Fux et al. (2024a); Bejan et al. (2024), dynamical complexity Leone et al. (2022); Garcia et al. (2023); Oliviero et al. (2024); Gu et al. (2024); Ahmadi and Greplova (2024); Odavic et al. (2024a), and (pseudo-)randomness Haferkamp et al. (2022a); Leone et al. (2021).
@@
Given the clear role each of these concepts individually play in both quantum information theory and many-body physics, one would hope to transitively understand the direct connection between the two. A question of foundational importance is therefore: How are the two disparate resources of magic and entanglement related to one other? The solution is not a priori obvious, at least in the usual quantum state setting. For instance, entanglement tends to grow maximally in Clifford circuits. On the other hand, product states can have high magic resource. And although details of the entanglement spectrum Chamon et al. (2014) can serve as witness to magic in a state Tirrito et al. (2024a); Turkeshi et al. (2023), even a single $T$-gate can make this spectrum consistent with random matrix theory predictions Zhou et al. (2020). Even more puzzlingly, there appears no clear connection to the integrability of the underlying dynamics in a many-body setting. Locally-interacting integrable and chaotic models alike produce linearly growing entanglement after a quench Calabrese and Cardy (2005); Chiara et al. (2006); Prosen and Znidarič (2007); Mezei and Stanford (2017); Nahum et al. (2017), while recent evidence suggests that state magic resources saturate in logarithmic time Turkeshi et al. (2024); Tirrito et al. (2024b); Odavic et al. (2024b). Despite this disparity, there has been recent interest in studying the interplay of, and boundary between, these fundamental resources Masot-Llima and Garcia-Saez (2024); Fux et al. (2024b); Gu et al. (2024); Frau et al. (2024); Bu et al. (2024).
@@
Figure 1: Depiction of key quantities investigated in this work. (a) The local-operator entanglement (LOE) is the entanglement $E_{A}^{(\alpha)}$ of the Choi state of an operator, $|O_{U}\rangle\!\rangle$, across some (doubled-space) spatial bipartition $A:\bar{A}$. (b) The extensive scaling of this quantity witnesses chaos, both in the sense of information scrambling Dowling et al. (2023) and non-integrability Prosen and Znidarič (2007). (c) In addition to being sensitive to entanglement, we show that the LOE equally probes magic resources and operator stabilizer entropies, unifying the three corresponding simulation techniques under the one umbrella.
@@
Here, we pose a new understanding of this relationship. We show that entanglement and magic are conceptually and quantitatively connected. We are able to do this by a shift in perspective from states to operators, noticing that free stabilizer resources are product in operator space. More specifically, consider an operator $O$ in the Heisenberg picture with respect to some evolution $U$, i.e., $O_{U}=U^{\dagger}OU$. By the Choi–Jamiołkowski isomorphism (CJI) $O_{U}$ has a dual pure state, $|O_{U}\rangle\!\rangle:=(O_{U}\otimes\mathds{1})\ket{\phi^{+}}$, where |$\ket{\phi^{+}}$ is the (normalized) maximally entangled state on a doubled Hilbert space. Entanglement and magic resources then have ready generalizations to operators through $|O_{U}\rangle\!\rangle$. The entanglement $E_{A}^{(\alpha)}(O_{U})$ of this state—once appropriately arranged into spatial partitions—is termed the local-operator entanglement (LOE) Prosen and Pižorn (2007), and has been studied in a variety of many-body systems Prosen and Znidarič (2007); Prosen and Pižorn (2007); Pižorn and Prosen (2009); Dubail (2017); Jonay et al. (2018); Alba et al. (2019); Bertini et al. (2020a, b); Alba (2021); see Fig. 1.
@@
The above relation Eq. (1) is the subject of the remainder of this work, with the full technical version to be found towards the end of the manuscript (Thms. 12). Our result sits in stark contrast to the corresponding state resource theories, where no such relation exists: states with high magic can be unentangled, while zero-magic (stabilizer) states can have maximal entanglement. Beyond the foundational significance of Eq. (1), there are operational implications. Chaotic dynamical systems have linearly scaling LOE with time, whereas it grows at-fastest logarithmically for integrable dynamics Prosen and Znidarič (2007); Prosen and Pižorn (2007); Pižorn and Prosen (2009); Dubail (2017); Jonay et al. (2018); Alba et al. (2019); Bertini et al. (2020a, b); Alba (2021). Our main result Eq. (1) additionally means that an extensive growth of magic resources is necessary ingredient for quantum chaos (see Cor. 3). Interpreting this result, chaotic growth of LOE implies the impossibility of efficient simulation of such dynamics using: tensor network, stabilizer, and Pauli truncation methods (see Cor. 4). LOE therefore serves as a unifying signature of non-classicality.
@@
Average Operator Entanglement from Magic.— Before detailing the full version of Eq. (1) (Thms. 12), we first review the pertinent monotones (summarized in Table 1) and introduce the relevant unitary ensembles (depicted in Fig. 2). Our many-body system of interest $\mathcal{H}$ is an $N$-qubit Hilbert space with total dimension $D=2^{N}$, and we denote by $\mathcal{U}_{N}$, $\mathcal{C}_{N}$, and $\mathcal{P}_{N}$ the $N$-qubit unitary, Clifford and Pauli groups respectively. By $U\sim\mathcal{U}_{N}$, we mean that $U$ is sampled uniformly according to the unitarily invariant (Haar) measure on $\mathcal{U}_{N}$, with an equivalent expression for the Cliffords.
@@
Consider a Heisenberg operator $O_{U}=U^{\dagger}OU$ for some non-trivial initial Pauli operator $O\in\mathcal{P}_{N}\backslash\{\mathds{1}\}$ and a unitary propagator $U$. Through the CJI, the LOE is defined as the bipartite entanglement of the Choi state $|O_{U}\rangle\!\rangle$ across a chosen bipartition $\mathcal{H}=\mathcal{H}_{A}\otimes\mathcal{H}_{\bar{A}}$, $E^{(\alpha)}_{A}(O_{U}):=S^{(\alpha)}(\operatorname{tr}_{\bar{A}}[|O_{U}\rangle\!\rangle\!\langle\!\langle O_{U}|])$, where $S^{(\alpha)}$ represents the (quantum) $\alpha$-Rényi entropy, $S^{(\alpha)}(\rho):=({1-\alpha})^{-1}\log(\operatorname{tr}[\rho^{\alpha}])$. Through abuse of notation, $\operatorname{tr}_{\bar{A}}$ refers to a partial trace over subsystem ${\bar{A}}$ in the doubled space $\mathcal{H}_{\bar{A}}\otimes\mathcal{H}_{\bar{A}}$ with dimension $D_{\bar{A}}^{2}$. Moreover, the operator purity of $O_{U}$ is defined in the usual way as $E^{(\text{pur})}_{A}(O_{U}):=\exp(-E^{(2)}_{A}(O_{U}))$. To first provide intuition on the interplay between LOE and magic resources, consider Clifford evolution, $C\in\mathcal{C}_{N}$. It is immediate to see from the definition of the Clifford group that starting from any initial Pauli operator, $O$ will evolve to some other Pauli operator and so the LOE is preserved to be zero. We again stress that no such relation between entanglement and magic holds for states. This invites the question of whether a more quantitative relation can be derived: if a unitary has only a few non-Clifford gates, how does the LOE grow? We will address this first by computing the average LOE for ensembles with a tunable magic monotone and show that there is a precise dependence.
@@
Figure 2: Any unitary can be decomposed as a global Clifford unitaries, $C_{0},C_{1}\in\mathcal{C}_{N}$, sandwiching a non-Clifford unitary on only $\nu(U)/2$ qubits, $V\in\mathcal{U}_{\nu}$ (top), where $\nu(U)$ is the unitary nullity Jiang and Wang (2023). A unitary with $T$-count $\tau(U)$ can be decomposed as layers of $N$-qubit Clifford unitaries $C_{i}\in\mathcal{C}_{N}$ interspersed with $\tau(U)$ single-site $T$-gates (bottom). The $\nu$-compressible ($\mu_{\nu}$) and $T$-doped ($\mu_{\tau}$) ensembles are constructed from uniformly sampling these components over the Clifford and unitary groups as appropriate.
@@
Consider deep Clifford circuits doped with some non-Clifford gates. That is, random Clifford circuits interspersed by $\tau(U)$ single-site $T$-gates, where $\tau(U)$ is the familiar $T$-count (see Table 1). If we randomize the Clifford components, such an ensemble of unitaries $\mu_{\tau}$ is called the $T$-doped Clifford ensemble Haferkamp et al. (2022a); Leone et al. (2021), as depicted in Fig. 2. Harnessing Weingarten techniques for averaging over the Clifford group Zhu et al. (2016); Roth et al. (2018); Leone et al. (2021), we can readily compute the operator purity for this example (see App.E), finding on-average that $\int_{U\sim\mu_{\tau}}{E^{(\text{pur})}_{A}(O_{U})}=\mathcal{O}(\exp(-\tau))$. Such a relationship points towards a proportionality between LOE and magic resources. However, while useful practically, the $T$-count of a unitary is not always well-defined.
A more concrete magic resource stemming from the algebraic structure of the Clifford group, is the stabilizer nullity Beverland et al. (2020). For quantum states |$\ket{\psi}$, this is defined in terms of the cardinality of its stabilizer group: the number of Pauli operators $P\in\mathcal{P}_{N}$ which satisfy $|\langle\psi|P|\psi\rangle|=1$. The nullity of a unitary operator can be defined analogously in terms of the Choi state $|U\rangle\!\rangle$ Jiang and Wang (2023),
$\nu(U):=2N-\log_{2}(|\mathrm{Stab}(\ket{U})|),$ (2)
with $\mathrm{Stab}(|U\rangle\!\rangle)=\{P_{1},P_{2}\in\mathcal{P}_{N}:\operatorname{tr}[P_{1}U^{\dagger}P_{2}U]/D=\pm 1\}$. The unitary nullity is a non-negative integer and upper-bounds the largest state nullity of $U$ acting on a stabilizer state. Moreover, it gives rise to a powerful representation theorem Leone et al. (2024a, b); Gu et al. (2024): that any unitary can be decomposed as
$U=C_{0}(V_{\ell}\otimes\mathds{1}_{N-\ell})C_{1},$ (3)
where $C_{0},C_{1}\in\mathcal{C}_{N}$ may act globally, while $V_{\ell}\in\mathcal{U}_{\ell}$ acts on $\nu/2\leq\ell\leq\nu$ qubits. Here, the lower-bound comes from a counting argument of the unitary stabilizer group generators, while the upper bound is shown explicitly in Ref. Leone et al. (2024b).
Returning to the question of the LOE’s sensitivity to magic resources, we can also compute the average-case operator purity for a unitary with a given unitary nullity $\nu(U)$. We define the $\nu$-compressible ensemble of unitaries $\mu_{\nu}$ as that generated from independently and uniformly sampling $C_{0},C_{1}\in\mathcal{C}_{N}$ and $V\in\mathcal{U}_{\ell}$ in Eq. (3), parametrized by an integer $\ell$. Here, for the Haar random sampling of $V$, the unitary nullity of $U$ will almost surely have nullity $\nu(U)=2\ell$. This comes from the invariance of unitary nullity under Clifford composition Jiang and Wang (2023), and from the fact that Haar random unitaries typically have maximal nullity Gu et al. (2024); Dowling et al. (2024):
$\nu(U)=\nu(V_{\ell}\otimes\mathds{1}_{N-\ell})=\nu(V_{\ell})\approx 2\ell.$ (4)
In the above, we have also used that nullity is additive under tensor products. Elements of $\mu_{\nu}$ are therefore typically defined by the decomposition $U\sim\mu_{\nu}$: $U=C_{0}(V_{\lceil\nu/2\rceil}\otimes\mathds{1}_{N-\lceil\nu/2\rceil})C_{1}$, and so we identify $\ell=\nu/2$ in the remainder of this work; see Fig. 2.
Similar to the case of the $T$-doped ensemble, from exact Weingarten techniques in App. D we show that operator purity on-average scales as $\int_{U\sim\mu_{\nu}}{E^{(\text{pur})}_{A}(O_{U})}=\mathcal{O}(\exp(-\nu))$.
@@
$T$-count Minimum number $\tau(U)$ such that $U=C_{0}\prod_{i=1}^{\tau}(TC_{i})$, for $C_{i}\in\mathcal{C}_{N}$.
@@
To clarify the relation of LOE and magic in the two discussed ensembles, we summarize the previous discussion in the following lower-bounds.
Theorem 1.
For Rényi indices of $\alpha\leq 2$, and $N_{A}=N/2$, the average LOE is bounded as
$\int_{U\sim\mathcal{E}}E^{(\alpha)}_{A}(O_{U})\geq\begin{cases}\nu-2-\mathcal{O}\left(\frac{2^{\nu}}{D}+\frac{1}{4^{\nu}}\right),&\mathcal{E}=\mu_{\nu},\\ \log\!\left(\frac{4}{3}\right)\tau-\mathcal{O}\left(\frac{(4/3)^{\tau}}{D}\right),&\mathcal{E}=\mu_{\tau}.\end{cases}$ (5)
Here, the lower-bounds are always positive, and $\mu_{\tau}$ and $\mu_{\nu}$ refer to the $T$-doped Clifford and $\nu$-compressible ensembles respectively.
@@
We first directly compute the LOE purity using (Clifford/unitary) Weingarten techniques in four-replica space Collins and Sniady (2006); Zhu et al. (2016); Roth et al. (2018); Leone et al. (2021), which led to the previous scalings discussed in the text. Then, after applying Jensen’s inequality for the negative logarithm involved in relating purity to $2-$Rényi entropy, together with the hierarchy of Rényi entropies ($S^{(\alpha)}\leq S^{(\beta)}$ for $\alpha\geq\beta$), we arrive directly at Eq. (5); see the Apps. B-E for details. ∎
In this result, we see a leading-order linear dependence of LOE with two magic monotones; c.f. Eq. (1). Interestingly, the dependence on $(3/4)^{\tau}$ in the $T$-count bound is also observed to leading order for the linear stabilizer Rényi entropy of the $T-$doped ensemble Leone et al. (2022); Haug et al. (2024).
Before moving on to an exact bound between LOE and magic resources (without averaging), we would like to compare the two lower-bounds of Eq. (5). The ensembles $\mu_{\tau}$ and $\mu_{\nu}$ coincide only for two cases: $\{\tau=\nu=0\}$, in which case they correspond uniform measure over Clifford group, and $\{\tau\to\infty,\nu\to 2N\}$, in which case they correspond to the unitary Haar ensemble. In the former case, as previously discussed, the LOE is trivially equal to zero. In the latter case, we can use our previous results to arrive at the Page-curve for the average LOE purity from Haar random dynamics,
@@
Note that this result is exact, improving upon a leading-order expression given in Ref. Kudler-Flam et al. (2021).
Unifying Entanglement and Magic.— Having studied the average case, we will now show that exact bounds can be derived between these two quantities for arbitrary Rényi indices and any given evolution, resulting in Eq. (1). In order to prove our main result, we first review one further magic monotone, operator stabilizer entropy (OSE). The OSE is defined as the entropy of the distribution of squared amplitudes of the Heisenberg operator $O_{U}$ written in the Pauli basis Dowling et al. (2024), $M^{(\alpha)}(O_{U}):=(1-\alpha)^{-1}\log\sum_{P\in\mathcal{P}}\left(D^{-1}\operatorname{tr}[O_{U}P]\right)^{2\alpha}$. This quantity generalizes a popular state measure of magic Leone et al. (2022); Haug and Piroli (2023); Leone and Bittel (2024), lower-bounding other magic monotones while satisfying a light-cone bound for local dynamics an dictating the efficiency of Pauli-truncation methods. Using the properties of OSE and unitary nullity, we are now in a position to re-express Eq. (1) in full technical detail.